# Revealing the $\chi_{\rm eff}$-$q$ correlation among Coalescing Binary Black Holes and Evidence for AGN-driven Hierarchical Mergers

### v1

@@ Appendix C Additional results
Figure 4 and Figure 5 present the full hyperparameters (except for the interpolation knots) of the Transition model, and Mixture model respectively. We find that whether we fix $a=0,b=0$ or not does not affect the transition / classification of the two subpopulations. These results are consistent with the analysis using spin-magnitude distributions in our previous work (Li et al., 2024b). We observe that the results of the two models are well consistent with each other. The first $\chi_{\rm eff}$ distribution narrowly peaks at $\sim 0.05$, while the second $\chi_{\rm eff}$ distribution peaks at $\sim 0.35$, favoring a contribution from hierarchical mergers in AGN disks (Yang et al., 2019).
@@
Appendix D Is the second $\chi_{\rm eff}$ distribution symmetric?
@@
The symmetry of $\chi_{\rm eff}$ distribution is critical for determining the formation environments of hierarchical mergers. Hierarchical mergers in star clusters always exhibit a symmetric $\chi_{\rm eff}$ distribution (Payne et al., 2024; Fishbach et al., 2022). While AGN-driven hierarchical mergers tend to favor positive $\chi_{\rm eff}$ due to gas torques (McKernan et al., 2018). In this section, we infer with variable lower edge and upper edge ($\chi_{\rm min,2}$, $\chi_{\rm max,2}$) for the second $\chi_{\rm eff}$ distribution. We infer that $\chi_{\rm max,2}>0.4$ at 97% credible level. Although, we cannot rule out the case of $\chi_{\rm min,2}=0$, it is more likely that $\chi_{\rm min,2}$ takes on a negative value, as shown in Figure 6.
@@
Figure 6: $\chi_{\rm eff}$ distributions inferred using Transition with variable edges $\chi_{\rm min,2}$ and $\chi_{\rm max,2}$ truncated on the second $\chi_{\rm eff}$ distribution.
Appendix E Are there $\chi_{\rm eff}-q$ correlation in subpopulations?
As presented in Section 3, we find no evidence for $\chi_{\rm eff}-q$ correlation in the low-spin (or first-generation) subpopulation. However, the first-generation subpopulation is most likely a mix of BBHs originating from both isolated and dynamical channels (Wang et al., 2022; Li et al., 2024a; Ray et al., 2024). Therefore, it is possible that the $\chi_{\rm eff}-q$ correlation occurs in a single, specific channel, such as the isolated evolution channel (Banerjee & Olejak, 2024).
We first artificially divide the entire population into three subpopulations according to our previous works (see Supplemental Table IV of Li et al., 2024b) and (Table 5 of Li et al., 2024a). The first is the low-spin & nearly-aligned subpopulation, that is related to the isolated evolution channels, see Table 4. The second is the high-spin subpopulation, related to hierarchical mergers, which are also identified within this work. The third is the low-spin & isotropic subpopulation that is associated with the first-generation BBHs formed in star clusters. The reason why we do not jointly fit the three subpopulation in this work is that, it is difficult to distinguish the first and third subpopulations using the $\chi_{\rm eff}$ distribution, since the effective spins in both subpopulations are close to zero as shown in Figure 3.
We then fit the three sets of samples with the Base (PS) model, respectively. Figure 7 shows the posterior distributions of the parameters describing the $\chi_{\rm eff}$ and $q$ distribution in each potential subpopulation. We find there is no evidence of correlation in either subpopulation, see also Table 3.
Interestingly, if we fit the combination of the second and third sub-set samples, which is potentially associated with the dynamical assembly (Li et al., 2024a), then we find strong evidence for the $\chi_{\rm eff}$-$q$ correlation with $\ln\mathcal{B}=2.6$, see also Figure 7 (Right). This result is consistent with the predictions of mergers in AGN disks (Cook et al., 2024). We find that a significant fraction of BBHs in the 2nd pop have large (positive) $\chi_{\rm eff}$ values, while the BBHs in the 3rd pop are located at $\chi_{\rm eff}\sim 0$,$q\sim 1$. Therefore, the $\chi_{\rm eff}$-$q$ anti-correlation likely arises from the superposition of these two types of subpopulations.
Figure 7: Left: posterior distribution of parameters describing the $\chi_{\rm eff}$-$q$ distribution in each subpopulation. Right: posterior $\chi_{\rm eff}$-$q$ distribution of individual event reweighed to the population-informed prior. The shaded areas mark the 90% credible regions and the points stand for the mean values.
Table 3: Model comparison
Data Sets $\ln{\mathcal{B}^{\rm with~{}corr.}_{\rm w./o.~{}corr.}}$
Whole -0.8
Low-Spin Aligned (Field; 1st pop) -1.2
High-Spin (2G; 2nd pop) -0.1
Low-Spin Isotropic (Dynamical 1G; 3rd pop) -1.0
Low-Spin (1G; 1st+3rd pop) -2.1
Low-Spin Isotropic + High-Spin (Dynamical; 2nd+3rd pop) 2.6
Note: Bayes factor for each potential subpopulation with versus without a $\chi_{\rm eff}-q$ correlation.
Table 4: Events of each potential subpopulation
Low-Spin Aligned (Field; 1st pop)
GW151012_095443 GW151226_033853 GW170608_020116 GW190412_053044 GW190512_180714 GW190707_093326
GW190708_232457 GW190720_000836 GW190728_064510 GW190828_065509 GW190924_021846 GW190930_133541
GW190725_174728 GW191105_143521 GW191129_134029 GW191204_171526 GW191216_213338 GW200202_154313
GW200316_215756 GW191103_012549
High-Spin (2G; 2nd pop)
GW170729_185629 GW190517_055101 GW190519_153544 GW190521_030229 GW190602_175927 GW190620_030421
GW190701_203306 GW190706_222641 GW190929_012149 GW190805_211137 GW191109_010717
Low-Spin Isotropic (Dynamical 1G; 3rd pop)
GW150914_095045 GW170104_101158 GW170809_082821 GW170814_103043 GW170818_022509 GW170823_131358
GW190408_181802 GW190413_134308 GW190421_213856 GW190503_185404 GW190513_205428 GW190521_074359
GW190527_092055 GW190630_185205 GW190727_060333 GW190803_022701 GW190828_063405 GW190910_112807
GW190915_235702 GW190925_232845 GW190413_052954 GW190719_215514 GW190731_140936 GW191127_050227
GW191215_223052 GW191222_033537 GW191230_180458 GW200112_155838 GW200128_022011 GW200129_065458
GW200208_130117 GW200209_085452 GW200219_094415 GW200224_222234 GW200225_060421 GW200302_015811
GW200311_115853 GW200216_220804
Note: The 2nd sub-set of GW events are figured out according to Li et al. (2024b), each event has probability $>0.5$ to
containing at least one high-spin BH. The 3rd sub-set are figured out according to Li et al. (2024a), each event has
probability $>0.5$ belonging to isotropic-spin sub-population.

### v2

@@ Appendix C Additional results
Figure 3 and Figure 4 present the full hyperparameters (except for the interpolation knots) of the Transition model, and Mixture model respectively. We find that whether we fix $a=0,b=0$ or not does not affect the transition / classification of the two subpopulations. These results are consistent with the analysis using spin-magnitude distributions in our previous work (Li et al., 2024b). We observe that the results of the two models are well consistent with each other. The first $\chi_{\rm eff}$ distribution narrowly peaks at $\sim 0.05$, while the second $\chi_{\rm eff}$ distribution peaks at $\sim 0.4$, favoring a contribution from hierarchical mergers in AGN disks (Yang et al., 2019). Note that some other formation channels, such as stable mass transfer (Bavera et al., 2021) and chemically homogeneous evolution (Mandel & de Mink, 2016) can also produce the $\chi_{\rm eff}$ distributions peaking at positive values.
@@
For the Mixture models, the maximum mass of the first subpopulation is not well constrained but exhibits a tail extending to larger values, see Figure 4. However, the masses of 99.5th and 99.75th, and 99.9th percentiles for the first subpopulation are well measured. Especially the $m_{99\%}$ of the first subpopulation differs significantly from that of the entire population inferred from the base models, see Figure 5.
Figure 6 presents the mass distributions of the subpopulations and the entire population inferred with different models. A consistent feature across all models (Mixture and Transition) is the rapid decline of the first subpopulation’s mass distribution above $\sim 40M_{\odot}$, aligning with the BH mass gap predicted by the Pair-instability Supernova (Farmer et al., 2019). The second subpopulation dominates the high-mass range; however, the lower boundary of its mass distribution remains uncertain. We additionally tested a more flexible mass function for the second subpopulation in the Mixture (PS) model by applying low-mass tapering and perturbation functions to the second power-law component (labeled PS+PS). Despite this modification, the results are nearly identical to those without low-mass tapering and perturbation functions, as demonstrated in Figure 6.
Figure 5: Posteriors of the maximum mass, masses of 99.5th and 99.75th, and 99.9th percentiles of the first subpopulation for the Mixture models. The contours mark the central 50% and 90% posterior credible regions, the values represent the median and 90% credible intervals.
Figure 6: Mass distributions inferred with various models. The thick and thin lines represent the mean values and 90% credible intervals.
Appendix D The second subpopulation
D.1 Is the second $\chi_{\rm eff}$ distribution symmetric?
@@
The symmetry of $\chi_{\rm eff}$ distribution is critical for determining the formation environments of hierarchical mergers. Hierarchical mergers in star clusters always exhibit a symmetric $\chi_{\rm eff}$ distribution about zero (Payne et al., 2024; Fishbach et al., 2022). While AGN-driven hierarchical mergers tend to favor positive $\chi_{\rm eff}$ due to gas torques (McKernan et al., 2018). In this section, we use two extended models to investigate the $\chi_{\rm eff}$ distribution of the second subpopulation. The accompanied mass function is the PS model. First, we infer with variable lower edge and upper edge ($\chi_{\rm min,2}$, $\chi_{\rm max,2}$) for the second $\chi_{\rm eff}$ distribution (labeled Truncated). We infer that $\chi_{\rm max,2}>0.4$ at 97% credible level. However, the $\chi_{\rm min,2}$ is only poorly constrained. Second, we implement a spline-based model for the second $\chi_{\rm eff}$ distribution (labeled Spline), $P_{\mathcal{S}}(\chi_{\rm eff})\propto e^{f(\chi_{\rm eff})}[-1,1]$. $f(x)$ is a cubic spline function defined by 6 nodes linearly located in $[-1,1]$. We restrict the 1-st and 6-th nodes to be -10 corresponding to $P_{\mathcal{S}}(\chi_{\rm eff}=1),P_{\mathcal{S}}(\chi_{\rm eff}=-1)\sim 0$, and the priors on the amplitude of 2-nd to 5-th nodes are unit Gaussian distributions (Golomb & Talbot, 2023).
Figure 7 presents the $\chi_{\rm eff}$ distributions inferred from the Truncated model and Spline model. The $\chi_{\rm eff}$ distributions of the first subpopulation show high consistency between both models. For the second subpopulation, both models exhibit a preference for positive $\chi_{\rm eff}$ values, though negative values cannot be ruled out based on the current constraints.
@@
Figure 7: $\chi_{\rm eff}$ distributions inferred using Transition (PS) with Truncated model and Spline model describing the second $\chi_{\rm eff}$ distribution. The thick and thin lines represent the mean values and 90% credible intervals.
D.2 Is there $\chi_{\rm eff}-q$ correlation in the second subpopulation?
To test whether there is $\chi_{\rm eff}-q$ correlation in the second subpopulation, we extend the Transition model, i.e., replacing the second $\chi_{\rm eff}$ distribution $\mathcal{G}_{[-1,1]}(\chi_{\rm eff}|\mu_{\chi,2},\sigma_{\chi,2})$ with $\mathcal{G}_{[-1,1]}(\mu_{\chi}(q;\mu_{\chi,2},a_{2}),\sigma_{\chi}(q;\sigma_{\chi,2},b_{2}))$. We find there is no evidence for the $\chi_{\rm eff}-q$ correlation in the second subpopulation, as shown in Figure 8. Therefore, comparing the Base models and the Transition / Mixture models, we can conclude that the $\chi_{\rm eff}-q$ anti-correlation in the entire population is mainly attributed to the superposition of two subpopulations.
Figure 8: Hyperparameters describing the $\chi_{\rm eff}-q$ correlation for the second subpopulation. The contours mark the central 50% and 90% posterior credible regions, the values represent the median and 90% credible intervals.
Appendix E Mock data studies
To test the validity of our findings, we conduct an end-to-end injection study, including the parameter estimation and hierarchical analysis of mock signals. We generate three distinct mock populations. The first has a non-evolving $\chi_{\rm eff}$ distribution (hereafter NonEvo). The second has the same feature that inferred with the Transition model in real data (hereafter TF&Asym). The third is similar to the second one, but for the $\chi_{\rm eff}$ of high-mass BBHs exhibiting a symmetric distribution about zero (hereafter TF&Sym). Then we use the non-evolving model and Transition model to recover the features in the mock populations.
For all three mock populations, the mass distributions are generated by the PP model with $m_{\rm min}=5~{}M_{\odot}$, $m_{\rm max}=80~{}M_{\odot}$, $\delta_{\rm m}=5~{}M_{\odot}$, $\alpha=3.5$, $\beta=1$, $\mu_{\rm m}=35~{}M_{\odot}$, $\sigma_{\rm m}=3~{}M{\odot}$, $\lambda_{\rm peak}=0.02$, the merger rate density as a function of redshift follows Eq. (B6) with $\gamma=2.7$, $z_{\rm p}=2$, and $\kappa=4$. For the NonEvo case, the $\chi_{\rm eff}$ distribution follows a Gaussian distribution $\mathcal{G}_{[-1,1]}(\mu_{\chi},\sigma_{\chi})$ with $\mu_{\chi}=0.06$ and $\sigma_{\chi}=0.1$. For the TF&Asym case, the $\chi_{\rm eff}$ distribution exhibits mass dependence, i.e., the subpopulation with primary masses $<40~{}M_{\odot}$ follows a Gaussian distribution $\mathcal{G}_{[-1,1]}(\mu_{\chi,1},\sigma_{\chi,1})$ with $\mu_{\chi,1}=0.05$ and $\sigma_{\chi,1}=0.05$, and the subpopulation with primary masses $>40~{}M{\odot}$ follows a Gaussian distribution $\mathcal{G}_{[-1,1]}(\mu_{\chi,2},\sigma_{\chi,2})$ with $\mu_{\chi,2}=0.3$ and $\sigma_{\chi,2}=0.3$. The TF&Sym case modifies the TF&Asym scenario by setting $\mu_{\chi,2}=0$ while maintaining other parameters identical..
The mock detections are sampled from the injection campaign for O3 Search Sensitivity Estimates [^4] https://doi.org/10.5281/zenodo.7890437 with the inverse FAR $>$ 1 yr and the weight proportional to $p(\theta|{\bf\Lambda})/p_{\rm draw}(\theta)$, where $p(\theta|{\bf\Lambda})$ is probability distribution of the mock population, and $p_{\rm draw}(\theta)$ is the probability distribution from which the injection campaigns are drawn. For each case, we adopt 69 events. Following Callister et al. (2021), we set aligned spin components to $\chi_{1,z}=\chi_{2,z}=\chi_{\rm eff}$ and in-plane components to zero. Then we use IMRPhenomXAS waveform(Pratten et al., 2020) to generate GW signal and inject it to the noise generated by the “O3 actual” noise power spectral densities [^5] https://dcc.ligo.org/LIGO-T2000012/public. We perform parameter estimation on each event using BILBY (Ashton et al., 2019), with the NESSAI nested sampler (Williams et al., 2025).
For each case, we employ the non-evolution model and transition model to recover the mock population, i.e., Base (PP) with $a,~{}b=0$ and Transition (PP) with $a,~{}b=0$. The recovered $\chi_{\rm eff}$ distributions are presented in Figure 9 (Top). It is shown that two distinct sub-populations are successfully recognized by the transition model in the TF&Sym and the TF&Asym mock populations. As anticipated, the second sub-population is ambiguous for the NonEvo mock population. For both the TF&Sym and the TF&Asym cases, the Transition model is more favored than the Base by $\ln\mathcal{B}\sim 8$. conversely, in the NonEvo case, the Transition model is less favored by $\ln\mathcal{B}\sim-1.5$.
To further assess the robustness of our analysis, we repeated the entire simulation process many times, including mock population generation, injection, recovery, and hierarchical inference. All runs produced consistent results, as illustrated in the bottom panel of Figure 9.
Figure 9: Top: $\chi_{\rm eff}$ distributions recovered from mock population with Transition model and non-evolving model, the black lines are for the injections. Bottom: the $\chi_{\rm eff}$ distributions of the second sub-population recovered from several datasets, the black lines are for the distribution of the second sub-populations. The thick and thin lines are for the mean values and 90% credible intervals.